# Criticality Leveraged Adversarial Training (CLAT) for Boosted Performance via Parameter Efficiency

### v2

@@ Abstract
Adversarial training enhances neural network robustness but suffers from a tendency to overfit and increased generalization errors on clean data. This work introduces CLAT, an innovative approach that mitigates adversarial overfitting by introducing parameter efficiency into the adversarial training process, improving both clean accuracy and adversarial robustness. Instead of tuning the entire model, CLAT identifies and fine-tunes robustness-critical layers—those predominantly learning non-robust features—while freezing the remaining model to enhance robustness. It employs dynamic critical layer selection to adapt to changes in layer criticality throughout the fine-tuning process. Empirically, CLAT can be applied on top of existing adversarial training methods, significantly reduces the number of trainable parameters by approximately 95%, and achieves more than a 2% improvement in adversarial robustness compared to baseline methods.
@@ 1 Introduction
Advancements in deep learning models have markedly improved image classification accuracy. Despite this, their vulnerability to adversarial attacks — subtle modifications to input images that mislead the model — remains a significant concern [9, 28]. The research community has been rigorously exploring theories to comprehend the mechanics behind adversarial attacks [4]. Ilyas et al. [14] uncover the coexistence of robust and non-robust features in standard datasets. Adversarial vulnerability largely stems from the presence of non-robust features in models trained on standard datasets, which, while highly predictive and beneficial for clean accuracy, are susceptible to noise [28]. Unfortunately, it is observed that deep learning models tend to preferentially learn these non-robust layers. Inkawhich et al. [15, 16] further demonstrate that adversarial images derived from the hidden features of certain intermediate non-robust/“critical” layers exhibit enhanced transferability to unseen models. This suggests a commonality in the non-robust features captured by these layers. While identifying these critical layers to improve their robustness is appealing, this process often requires the costly generation of attacks against each individual network layer. Efficient methods to identify and effectively address the criticality of such layers are still lacking.
In contrast to layer-wise feature vulnerability analysis, adversarial training [3, 19, 8], involves training entire neural networks with adversarial examples generated in real-time. This approach inherently encourages all layers in the model to learn robust features from adversarial images, thereby enhancing the model’s resilience against attacks. However, given the more challenging optimization process of learning from adversarial examples than from clean examples, adversarial training also brings hurdles such as heightened errors on clean data and susceptibility to overfitting, ultimately diminishing its effectiveness in practical applications [25, 36, 23, 17]. Despite various efforts to enhance adversarial training, such as modifying input data and adjusting loss functions [12, 23, 36, 30, 32, 21], they still frequently fall short in alleviating the aforementioned challenges.
@@
Figure 1: CLAT overview. CLAT fine-tunes the selected critical layers (red) while freezing other layers (grey). fine-tuning objective is computed per Equation 9. Critical layers are adjusted periodically. Pseudocode is provided in Appendix A.
@@
In light of these challenges, we introduce CLAT, a paradigm shift in adversarial training, where we mitigate overfitting during adversarial training by identifying and tuning only the robustness-critical model layers. CLAT commences by pinpointing critical layers within a model using our novel, theoretically grounded, and easily computable,“criticality index”, which we developed to identify layers which have learned non-robust features dominantly. Subsequently, our algorithm meticulously fine-tunes these critical layers to remove their non-robust features and reduce their criticality, while freezing the other, non-critical layers. Dynamic selection of critical layers is conducted during the training process to always focus the fine-tuning on the most-in-need layers, avoiding the overfitting of full-model adversarial training. CLAT therefore enhances both clean and adversarial accuracy compared to previous adversarial training methods. Besides reducing overfitting, the critical layer selection in training also makes CLAT a streamlined, parameter-efficient fine-tuning method, enabling memory savings with its significantly reduced trainable parameters.
@@
  • We introduce the “criticality index”, a quantitative metric designed to efficiently identify critical layers for the adversarial vulnerability of the model.
@@
  • We propose CLAT, a parameter-efficient adversarial fine-tuning algorithm focusing on layer criticality reduction. CLAT mitigates overfitting in the adversarial training process, and applies to various model training scenarios and baseline adversarial training methods.
CLAT stands out by markedly reducing overfitting risks, maintaining or even slightly improving the clean accuracy while significantly enhancing adversarial resilience by at least 2% with a 96% reduction in trainable parameters.
@@ 2.1 Adversarial training
Adversarial Training (AT) was first introduced by Goodfellow et al. [9], who demonstrated how the integration of adversarial examples into the training process could substantially improve model robustness. This idea evolved into a sophisticated minimax optimization approach with Projected Gradient Descent Adversarial Training (PGD-AT) [19], which employs PGD attacks in training. Regarded as the gold standard in AT, PGD-AT generates adversarial training samples using multiple steps of projected gradient descent, leading to substantially improved empirical robustness [6, 3, 8]. Further refining this approach, TRADES [36] optimizes a novel loss function to balance classification accuracy with adversarial robustness. Recent enhancements in AT, including model ensembling and data augmentation, have also produced notable improvements in model resilience. [33, 34, 7]Inkawhich et al. [15] propose “Activation Attacks” (AA) which underscore the efficacy of leveraging intermediate model layers for generating stronger adversarial attacks, suggesting that incorporating AA in adversarial training could fortify defenses. Their findings provide a foundation for our method which integrates these intermediate critical layers into our adversarial training strategy.
2.2 Efficient adversarial training
Adversarial training methods like PGD-AT and TRADES are computationally intensive, necessitating the multi-step generation of adversarial examples, optimization of complex objective functions, and comprehensive model tuning [26]. Numerous efforts to enhance the efficiency of adversarial training have been proposed. Shafahi et al. [26] introduced an efficient model known as “Free” AT, using a single backpropagation step for both training and PGD adversary generation to accelerate adversarial training. However, this approach struggled with maintaining model robustness and increased overfitting, primarily due to gradient alignment issues. Wong et al. [31] then proposed “Fast” AT, which aimed to address these shortcomings, but was later found to be vulnerable to similar problems [2]. As a result, Andriushchenko and Flammarion [2] developed GradAlign but this approach tripled the training time due to second-order gradient computation. Lastly, RiFT leverages network layer redundancies to improve general performance, but its heuristic-based layer redundancy measurement limits final robustness [37]. In contrast, CLAT uses a theoretically grounded, dynamic, critical layer selection mechanism, resulting in a more robust and parameter-efficient adversarial training algorithm. Furthermore, CLAT does not require a fully trained model and can be combined with existing fast-AT methods to enhance performance and mitigate overfitting.
@@ 3 Methods
Building on prior attack and defense research [15, 16, 37] which demonstrates that not all model layers equally learn non-robust features, we aim to improve model robustness and parameter efficiency by identifying and fine-tuning only those critical layers that are prone to learning non-robust features, while keeping the non-critical layers frozen. In this section, we begin by defining and identifying critical layers, then outline our objectives for reducing their criticality. Finally, we present our complete CLAT algorithm, which effectively enhances model robustness.
@@ 3.1 Layer Criticality
Under adversarial perturbation, features from all layers will be altered, leading to incorrect outputs. To understand how the learned parameters of the model affect the features under perturbation, we consider a function $G_{i}(\cdot)$ defined at the output of layer $i$ on inputs close to a clean data point $x$:
$G_{i}(z)=||F_{i}(z)-F_{i}(x)||_{2}^{2}.$ (2)
Clearly, $z=x$ is a minima of $G_{i}(z)$. The worst-case curvature of the function $G_{i}$ at the neighborhood of $z=x$ can be estimated following the formulation by Moosavi-Dezfooli et al. [20] as
where $x^{\prime}$ is a worst-case perturbation (adversarial attack) maximizing $G_{i}(z)$ in the vicinity of $x$, and $\nabla G_{i}(x)=0$ by definition given it is a minimum.
The numerator in the curvature formulation can be further derived as
Following the observation in [20], a higher curvature indicates the feature to be more sensitive to adversarial examples. We can therefore use the curvature formulation $\nu_{i}(x)$ under a fixed perturbation budget $||x^{\prime}-x||_{p}\leq\epsilon$ to estimate the layer sensitivity, or weakness.
In practice, it is difficult to explicitly instantiate $\frac{\partial F_{i}(x^{\prime})}{\partial x^{\prime}}$ for a neural network. To this end, we simplify the formulation in Equation 4 by assuming $\frac{\partial F_{i}(x^{\prime})}{\partial x^{\prime}}$ as a uniform vector. This leads to our definition of the $\epsilon$-weakness of layer $i$’s feature as:
@@
where $N_{i}$ denotes the dimensionality of the output features at layer $i$, therefore normalizing the weakness measurement of layers with different output sizes. The weakness measurement is proportional to the curvature estimation in Equation 4. A higher weakness value indicates that the feature vector is more vulnerable to input perturbations. The functionality of cascading layers from 1 to $i$ affects the vulnerability of the hidden features, as described by this formulation.
For the purpose of efficient fine-tuning, we want to identify the layers that are the most critical to the lack of robustness, characterized by their increased susceptibility to perturbations from adversarial inputs. We therefore provide the following definition:
@@
Critical layer: A layer is considered critical if it exhibits a greater propensity to learn non-robust features or demonstrates heightened sensitivity to adversarial input perturbations relative to other layers in the model.
@@
As a sanity check, the feature weakness at the output of layer $i$ can be attributed to the criticality of all previous layers as $\mathcal{W}_{\epsilon}(F_{i})=\prod_{k=1}^{i}\mathcal{C}_{f_{k}}$. Conversely, a layer with a larger criticality index will increase the weakness of the features after it, indicating the layer is critical according to Definition 3.1, as it heightens the feature sensitivity to the adversarial input.
One drawback of the formulation in Equation (6) is that computing the feature weakness involves finding the worst-case perturbation against the hidden features at each layer, which is a costly process to conduct sequentially for all layers. In practice, we find it possible to approximate the worst-case perturbation against features with an untargeted PGD attack against the model output, so that we can use the same PGD perturbation $\delta$ to estimate the feature weakness of all layers following Equation 5. In this way, with a reasonably sufficient batch size, we can compute the critical indices for all layers in a model with two forward passes: one with the clean input $x$ and one with the PGD attack input $x+\delta$. We make the following proposition:
@@ 3.2 Criticality-targeted fine-tuning
Once the critical layers are identified, we fine-tune them to reduce their criticality, thereby decreasing the weakness of subsequent hidden features and enhancing model robustness. For a critical layer $i$, we optimize the trainable parameters to minimize $\mathcal{C}_{f_{i}}$. Note that in the criticality formulation in Equation 6, the weakness of the previous layer’s output, $\mathcal{W}_{\epsilon}(F_{i-1})$, is constant with respect to $f_{i}$. Thus, the optimization objective for $f_{i}$ can be simplified as
@@
where a single perturbation is utilized to capture the weakness across all critical layers. In practice, a projected gradient ascent optimization with random start is used to solve the inner maximization.
Minimizing the objective in Equation 8 by adjusting the trainable variables of the critical layers will reduce their feature weaknesses. However, the removal of non-robust features in these layers may affect the functionality of the model on clean inputs. As a tradeoff, we also include the cross entropy loss $\mathcal{L}(\cdot)$ in the final optimization objective, which derives the optimization objective on the critical layers during the fine-tuning process
@@
where the hyperparameter $\lambda$ serves as a balancing factor between the two loss terms. Note that only the selected critical layers $f_{S}$ are optimized in Equation 9 while the other non-critical layers are frozen, justifying our fine-tuning to be parameter-efficient.
@@ 3.3 CLAT adversarial training
Similar to parameter-efficient fine-tuning techniques in other domains, we design CLAT as a fine-tuning approach, which is applied to neural networks that have undergone preliminary training. The pretraining phase allows all layers in the model to capture useful features, which will facilitate the identification of critical layers in the model. Notably, CLAT’s versatility allows it to adapt to various types of pretrained models, either adversarially trained or trained on a clean dataset only. In practice, we find that models do not need to fully converge during the pretraining phase to benefit from CLAT fine-tuning. For example, in case of the CIFAR-10 dataset, 50 epochs of PGD-AT training would be adequate. We consider the number of adversarial pretraining epochs as a hyperparameter and provide further analysis on the impact of pretraining epochs in Section 4.3.1.
After the pretraining, CLAT begins by identifying and selecting critical layers in the pretrained model. We then fine-tune critical layers only while freezing the rest of the layers. This process is illustrated in Figure 1, and the pseudocode is provided in Algorithm 1 in Appendix A for greater clarity. Compared to full-model fine-tuning, CLAT not only improves parameter efficiency with fewer trainable parameters, but also reduces overfitting, a common issue in standard adversarial training. As fine-tuning progresses, the critical layers will be updated to reduce their criticality, making them less critical than some of the previously frozen layers. Subsequently, we perform periodic reevaluation of the top $k$ critical indices, ensuring continuous adaptation and optimization of the layers that are the most in need in the training process. Through hyperparameter optimization, we find 10 epochs to be adequate to optimize the selected critical layers for all models that we tested.
@@ 4.1 Experimental settings
We conducted experiments using two widely recognized image classification datasets, CIFAR10 and CIFAR100. Each dataset includes 60,000 color images, each 32×32 pixels, divided into 10 and 100 classes respectively [18]. For our experiments, we deployed a suite of network architectures: WideResnets (34-10, 70-16) [35], ResNets (50, 18) [11], DenseNet-121 [13], PreAct ResNet-18[10], and VGG-19 [27]. In this paper, these architectures are referred to as WRN34-10, WRN70-16, RN50, RN18, DN121, Preact RN18 and VGG19 respectively.
@@
Since CLAT can be layered over clean pretraining, partial training, or other adversarial methods, results incorporating CLAT are denoted in our tables as "X+ CLAT," where “X” specifies the baseline method used prior to applying CLAT. Typically, this baseline method was applied for the first 50 epochs, followed by a fine-tuning phase during which CLAT was run for an additional 50 epochs. Models trained exclusively with PGD-AT typically require 150 epochs to achieve reported performances. We use PGD for attack generation unless stated otherwise in the baseline method, which employs a random start [19], with an attack budget of $\epsilon$ = 0.03 under the $\ell_{\infty}$ norm, an update step size of $\alpha$ = 0.007, and 10 attack steps. Experiments were conducted on a Titan XP GPU, starting with an initial learning rate of 0.1, which was adjusted according to a cosine decay schedule. To ensure the reliability of robustness measurements, we conducted each experiment a minimum of 10 times, reporting the lowest adversarial accuracies we observed.
@@ 4.2 CLAT performance
Table 1 presents the clean accuracy and white-box $\ell_{\infty}$ PGD [6] accuracy across various adversarial training techniques. The versatility and effectiveness of CLAT are demonstrated by combining it with various standard adversarial training methods. CLAT mitigates the overfitting commonly observed in traditional adversarial training methods, improving both clean and adversarial accuracy compared to the corresponding baselines.
We also show that reducing trainable parameters alone does not guarantee a performance improvement by comparing CLAT against other parameter-efficient fine-tuning methods like LoRA [1] and RiFT [37]. Our theoretically-grounded criticality indices enable the precise identification of critical layers and the focused elimination of non-robust features from these pivotal layers. Furthermore, we verify that fast adversarial training techniques as in [31] can be applied to solve the inner maximization in the CLAT training objective in Equation 8, where the “CLAT (Fast)” method also improves performance and robustness over Fast-AT baselines.
Besides the main table, we provide additional robustness results in Appendix B, where we showcase CLAT’s improved robustness against a variant of the white-box Auto Attack [8] in LABEL:wb_aa and verify the generalizability of the robustness to attacks of various strengths in Figure 4.
@@
PREACT RN18 fast-at [31] 81.46 45.55 50.10 27.72
@@
Table 2 evaluates the robustness against black-box attacks between models trained solely using PGD-AT and those augmented with CLAT. Attack settings are the same as those of the white-box attacks. As a sanity check, the accuracies under black-box attack surpass those observed under white-box scenarios, indicating that gradient masking does not appear in the CLAT model, and that the white-box robustness evaluation is valid. More significantly, models trained with CLAT consistently outperform those trained with PGD-AT, maintaining superior resilience in both black-box and white-box settings, regardless of the attack method or models employed.

### v3

# Boosting Adversarial Robustness with CLAT: Criticality-Leveraged Adversarial Training

@@ Abstract
Adversarial training (AT) is a common technique for enhancing neural network robustness. Typically, AT updates all trainable parameters, but such comprehensive adjustments can lead to overfitting and increased generalization errors on clean data. Research suggests that fine-tuning specific parameters may be more effective; however, methods for identifying these essential parameters and establishing effective optimization objectives remain unclear and inadequately addressed. We present CLAT, an innovative adversarial fine-tuning algorithm that mitigates adversarial overfitting by integrating "criticality" into the training process. Instead of tuning the entire model, CLAT identifies and fine-tunes fewer parameters in robustness-critical layers—those predominantly learning non-robust features—while keeping the rest of the model fixed. Additionally, CLAT employs a dynamic layer selection process that adapts to changes in layer criticality during training. Empirical results demonstrate that CLAT can be seamlessly integrated with existing adversarial training methods, enhancing clean accuracy and adversarial robustness by over 2% compared to baseline approaches.
@@ 1 Introduction
Advancements in deep learning models have markedly improved image classification accuracy. Despite this, their vulnerability to adversarial attacks — subtle modifications to input images that mislead the model — remains a significant concern (Goodfellow et al., 2015; Szegedy et al., 2014). The research community has been rigorously exploring theories to comprehend the mechanics behind adversarial attacks (Bai et al., 2021). Ilyas et al. (2019) uncover the coexistence of robust and non-robust features in standard datasets. Adversarial vulnerability largely stems from the presence of non-robust features in models trained on standard datasets, which, while highly predictive and beneficial for clean accuracy, are susceptible to noise (Szegedy et al., 2014). Unfortunately, it is observed that deep learning models tend to preferentially learn these non-robust features. Inkawhich et al. (2019; 2020) further demonstrate that adversarial images derived from the hidden features of
@@
Figure 1: CLAT overview. CLAT fine-tunes the selected critical layers (red) while freezing other layers (grey). fine-tuning objective is computed per Eq. 6. Critical layers are adjusted periodically. Pseudocode is provided in Appendix A.
certain intermediate non-robust/“critical” layers exhibit enhanced transferability to unseen models. This suggests a commonality in the non-robust features captured by these layers. While identifying these critical layers to improve their robustness is appealing, this process often requires the time-consuming generation of attacks against each individual layer. Methods to identify and effectively address the criticality of such layers are still lacking.
In contrast to layer-wise feature vulnerability analysis, adversarial training (Athalye et al., 2018; Madry et al., 2019; Croce & Hein, 2020), involves training entire neural networks with adversarial examples generated in real-time. This approach inherently encourages all layers in the model to learn robust features from adversarial images, thereby enhancing the model’s resilience against attacks. However, given the more challenging optimization process of learning from adversarial examples than from clean ones, adversarial training also brings hurdles such as heightened errors on clean data and susceptibility to overfitting, ultimately diminishing its effectiveness in practical applications (Schmidt et al., 2018; Zhang et al., 2019; Raghunathan et al., 2019; Javanmard et al., 2020). Despite various efforts to enhance adversarial training, such as modifying input data and adjusting loss functions (Hitaj et al., 2021; Raghunathan et al., 2019; Zhang et al., 2019; Wang et al., 2020; Wu et al., 2020b; Pang et al., 2022), these approaches still frequently fall short in alleviating the aforementioned issues.
@@
In light of these challenges, we introduce CLAT, a paradigm shift in adversarial training, where we mitigate overfitting during adversarial training by identifying and tuning only the robustness-critical model layers. CLAT commences by pinpointing critical layers within a model using our novel, theoretically grounded, and easily computable,“criticality index”, which we developed to identify layers which have learned non-robust features dominantly. Subsequently, our algorithm meticulously fine-tunes these critical layers to remove their non-robust features and reduce their criticality, while freezing the other, non-critical layers. Dynamic selection of critical layers is conducted during the training process to always focus fine-tuning on the most-in-need layers, avoiding the overfitting of full-model adversarial training. CLAT therefore achieves both clean and adversarial state-of-the-art (SOTA) accuracy compared to previous adversarial training methods.
@@
  • We introduce the “criticality index”, a quantitative metric designed to identify critical layers for the adversarial vulnerability of a model with minimal overhead.
@@
  • We propose CLAT, an adversarial fine-tuning algorithm that mitigates overfitting by focusing on reducing the criticality of fewer than 4% of trainable parameters. CLAT integrates seamlessly into diverse model training scenarios and baseline adversarial training methods.
CLAT stands out by markedly reducing overfitting risks, boosting both clean accuracy and adversarial resilience by approximately 2%.
@@ 2.1 Adversarial training
Adversarial Training (AT) was first introduced by (Goodfellow et al., 2015), who demonstrated how the integration of adversarial examples into the training process could substantially improve model robustness. This idea evolved into a sophisticated minimax optimization approach with Projected Gradient Descent Adversarial Training (PGD-AT) (Madry et al., 2019), which employs PGD attacks in training. Regarded as the gold standard in AT, PGD-AT generates adversarial training samples using multiple steps of projected gradient descent, leading to substantially improved empirical robustness (Carlini & Wagner, 2017; Athalye et al., 2018; Croce & Hein, 2020). Further refining this approach, TRADES (Zhang et al., 2019) optimizes a novel loss function to balance classification accuracy with adversarial robustness. Recent enhancements in AT, including model ensemble and data augmentation, have also produced notable improvements in model resilience. (Xie et al., 2020; Yang et al., 2020; Carmon et al., 2022)Inkawhich et al. (2019) propose “Activation Attacks” (AA) which underscore the efficacy of leveraging intermediate model layers for generating stronger adversarial attacks, suggesting that incorporating AA in adversarial training could fortify defenses. Their findings provide a foundation for our method which integrates these intermediate critical layers into our adversarial training strategy.
2.2 Adversarial training improvements and robust overfitting
Adversarial training methods like PGD-AT and TRADES are computationally expensive and prone to overfitting, requiring multi-step adversary generation, complex objectives, and extensive model tuning (Shafahi et al., 2019). To improve efficiency, Shafahi et al. (2019) proposed “Free” AT, which accelerates training by using a single backpropagation step for both training and PGD adversary generation. However, gradient alignment issues led to reduced robustness and increased overfitting. Similarly, Wong et al. (2020) introduced “Fast” AT, but it also suffered from similar weaknesses (Andriushchenko & Flammarion, 2020), prompting the development of GradAlign. Unfortunately, GradAlign tripled training time due to second-order gradient computation. Later efforts sought to address robust overfitting. RiFT (Zhu et al., 2023) improved general performance by leveraging layer redundancies but was constrained by heuristic redundancy measurements. Zhang et al. (2024) mitigated overfitting by disentangling natural and adversarial objectives, yet model-wide adjustments still limited robustness. In contrast, CLAT uses a theoretically grounded, dynamic, critical layer selection mechanism, resulting in improved robust generalization by tuning a critical subset of layers. Furthermore, CLAT is agnostic to attack generation methods in the AT processing, making it an ideal complement to existing fast-AT methods to mitigate overfitting.
@@ 3 Methods
Building on prior attack and defense research (Inkawhich et al., 2019; 2020; Zhu et al., 2023) which demonstrates that not all model layers equally learn non-robust features and having all layers learn robust features leads to overfitting, we aim to improve model robustness by identifying and fine-tuning only those critical layers that are prone to learning non-robust features, while keeping the non-critical layers frozen. In this section, we begin by defining and identifying critical layers, then outline our objectives for reducing their criticality. Finally, we present our complete CLAT algorithm, which effectively mitigates overfitting and enhances model robustness.
@@ 3.1 Layer Criticality
Under adversarial perturbation, features from all layers will be altered, leading to incorrect outputs. Following previous work (Hein & Andriushchenko, 2017; Finlay et al., 2018), the robustness, or weakness, of the feature can be linked to the local Lipschitz constant of function $F_{i}(\cdot)$. For a easier computation, we consider the worst-case feature difference under a fixed input perturbation budget of $\epsilon$. This leads to our definition of the $\epsilon$-weakness of layer $i$’s feature as:
@@
where $N_{i}$ denotes the dimensionality of the output features at layer $i$, therefore normalizing the weakness measurement of layers with different output sizes. The weakness measurement is proportional to the local Lipschitz constant. A higher weakness value indicates that the feature vector is more vulnerable to input perturbations. The functionality of cascading layers from 1 to $i$ affects the vulnerability of the hidden features, as described by this formulation.
Alternatively, Moosavi-Dezfooli et al. (2019) suggests the local curvature can be a more accurate estimation of robustness than the Lipschitz constant. However, the computation and optimization of curvature involves costly higher-order gradient computation. We provide additional derivations in Appendix E to show the feature weakness defined in Eq. 2 is also an effective approximation to the local curvature value.
For the purpose of mitigating overfitting, we want to identify the layers that are the most critical to the lack of robustness, characterized by their increased susceptibility to perturbations from adversarial inputs. Other already-robust layers shall then be fixed to avoid further overfitting to the adversarial training objective. We therefore provide the following definition:
@@
Critical layer: A layer is considered critical if it exhibits a greater propensity to learn non-robust features or demonstrates diminished robustness to adversarial input perturbations relative to other layers in the model.
@@
As a sanity check, the feature weakness at the output of layer $i$ can be attributed to the criticality of all previous layers as $\mathcal{W}_{\epsilon}(F_{i})=\prod_{k=1}^{i}\mathcal{C}_{f_{k}}$. Conversely, a layer with a larger criticality index will increase the weakness of the features after it, indicating the layer is critical according to Definition 3.1, as it weakens the feature robustness to the adversarial input.
One drawback of the formulation in Eq. 3 is that computing the feature weakness involves finding the worst-case perturbation against the hidden features at each layer, which is a costly process to conduct sequentially for all layers. In practice, we approximate the worst-case perturbation against features with an untargeted PGD attack against the model output, so that we can use the same PGD perturbation $\delta$ to estimate the feature weakness of all layers following Eq. 2. In this way, with a reasonably sufficient batch size, we can compute the critical indices for all layers in a model with two forward passes: one with the clean input $x$ and one with the PGD attack input $x+\delta$. We make the following proposition:
@@ 3.2 Criticality-targeted fine-tuning
Once the critical layers are identified, we fine-tune them to reduce their criticality, thereby decreasing the weakness of subsequent hidden features and enhancing model robustness. For a critical layer $i$, we optimize the trainable parameters to minimize $\mathcal{C}_{f_{i}}$. Note that in the criticality formulation in Eq. 3, the weakness of the previous layer’s output, $\mathcal{W}_{\epsilon}(F_{i-1})$, is constant with respect to $f_{i}$. Thus, the optimization objective for $f_{i}$ can be simplified as
@@
where a single perturbation is utilized to capture the weakness across all critical layers. In the common setting, a projected gradient ascent optimization with random start is used for the inner maximization.
Minimizing the objective in Eq. 5 by adjusting the trainable variables of the critical layers will reduce their feature weaknesses. However, the removal of non-robust features in these layers may affect the functionality of the model on clean inputs. As a tradeoff, we also include the cross entropy loss $\mathcal{L}(\cdot)$ in the final optimization objective, which derives the optimization objective on the critical layers during the fine-tuning process
@@
where the hyperparameter $\lambda$ serves as a balancing factor between the two loss terms. Note that only the selected critical layers $f_{S}$ are optimized in Eq. 6 while the other non-critical layers are frozen, preventing them from further overfitting.
@@ 3.3 CLAT adversarial training
We design CLAT as a fine-tuning approach, which is applied to neural networks that have undergone preliminary training. The pretraining phase allows all layers in the model to capture useful features, which will facilitate the identification of critical layers in the model. Notably, CLAT’s versatility allows it to adapt to various types of pretrained models, either adversarially trained or trained on a clean dataset only. In practice, we find that models do not need to fully converge during the pretraining phase to benefit from CLAT fine-tuning. For example, in case of the CIFAR-10 dataset, 50 epochs of PGD-AT training would be adequate. We consider the number of adversarial pretraining epochs as a hyperparameter and provide further analysis on the impact of pretraining epochs in Section 4.3.1.
After the pretraining, CLAT begins by identifying and selecting critical layers in the pretrained model. We then fine-tune critical layers only while freezing the rest of the layers. This process is illustrated in Fig. 1, and the pseudocode is provided in Algorithm 1 in Appendix A for greater clarity. As fine-tuning progresses, the critical layers will be updated to reduce their criticality, making them less critical than some of the previously frozen layers. Subsequently, we perform periodic reevaluation of the top $k$ critical indices, ensuring continuous adaptation and optimization of the layers that are the most in need in the training process. Through hyperparameter optimization, we find 10 epochs to be adequate to optimize the selected critical layers for all models that we tested.
@@
We conducted experiments using two widely recognized image classification datasets, CIFAR10 and CIFAR100. Each dataset includes 60,000 color images, each 32×32 pixels, divided into 10 and 100 classes respectively (Krizhevsky & Hinton, 2009). For our experiments, we deployed a suite of network architectures: WideResnets (34-10, 70-16) (Zagoruyko & Komodakis, 2017), ResNets (50, 18) (He et al., 2016b), DenseNet-121 (Huang et al., 2017), PreAct ResNet-18(He et al., 2016a), and VGG-19 (Simonyan & Zisserman, 2015). In this paper, these architectures are referred to as WRN34-10, WRN70-16, RN50, RN18, DN121, Preact RN18 and VGG19 respectively.
@@
Since CLAT can be layered over clean pretraining, partial training, or other adversarial methods, results incorporating CLAT are denoted in our tables as "X + CLAT," where "X" refers to the baseline method applied prior to CLAT. Typically, this baseline method is run for the first 50 epochs, followed by fine-tuning during which CLAT is applied for an additional 50 epochs. The total number of epochs is in line with the 100 epochs used in previous PGD-based adversarial training work (Zhang et al., 2019; Zhu et al., 2023).
For our baseline, we use PGD for attack generation during training, following a random start (Madry et al., 2019), with an attack budget of $\epsilon=0.03$ under the $\ell_{\infty}$ norm, a step size of $\alpha=0.007$, and 10 attack steps. The same settings apply to PGD attack evaluations. AutoAttack evaluations (Croce & Hein, 2020) also use a budget of $\epsilon=0.03$ under the $\ell_{\infty}$ norm, with no restarts for untargeted APGD-CE, 9 target classes for APGD-DLR, 3 target classes for Targeted FAB, and 5000 queries for Square Attack. These settings remain consistent unless explicitly noted otherwise.
Experiments were conducted on a Titan XP GPU, starting with an initial learning rate of 0.1, which was adjusted according to a cosine decay schedule. To ensure the reliability of robustness measurements, we conducted each experiment a minimum of 10 times, reporting the lowest adversarial accuracies we observed.
@@
Table 1 and Table 2 present white-box evaluation results using the PGD and Auto attack frameworks, respectively. These tables illustrate CLAT’s versatility and effectiveness when combined with various standard adversarial training methods, including state-of-the-art benchmarks from RobustBench (Croce et al., 2020). CLAT effectively mitigates the overfitting typically observed in traditional adversarial training, thereby enhancing both clean and adversarial accuracy compared to baseline methods.
Table 1 further highlights that reducing trainable parameters alone does not necessarily lead to improved performance. CLAT surpasses other parameter-efficient fine-tuning methods, such as LoRA (Aleem et al., 2024) and RiFT (Zhu et al., 2023), due to its ability to precisely identify critical layers and eliminate non-robust features from these layers. Additionally, we show that fast adversarial training techniques, as discussed by Wong et al. (2020), can be applied to address the inner maximization problem in the CLAT training objective described in Eq. 5. The "CLAT (Fast)" method not only enhances performance but also improves robustness compared to Fast-AT baselines.
Notably, CLAT models are trained with PGD-like attacks on hidden features without seeing Auto Attacks directly, but their robustness persists under these attacks (see Table 2). This suggests that different attacks across various networks share similarities in exploiting non-robust features. By addressing these non-robust features through critical layer fine-tuning, CLAT ensures that the robustness is adaptable across different attack settings and models. Lastly, we assess the robustness across various attack strengths in Appendix B.
@@
autolora (Zhang et al., 2024) 84.2 54.27 62.1 32.71
autolora + clat 86.45 58.53 64.91 35.82
PREACT RN18 fast-at (Wong et al., 2020) 81.46 45.55 50.10 27.72
@@
Table 2: Adversarial Accuracy on CIFAR-10 and CIFAR-100 when subjected to AutoAttack (AA).
Network Method Adversarial Accuracy (%)
CIFAR-10 CIFAR-100
DN121 pgd-at 47.56 23.13
pgd-at + clat 49.91 25.74
WRN70-16 pgd-at 54.32 28.25
pgd-at + clat 57.64 30.98
(Carlini & Wagner, 2017) 66.1 -
(Carlini & Wagner, 2017) + clat 68.43 -
RN50 pgd-at 46.22 23.48
pgd-at + clat 49.45 25.81
(Madry et al., 2019) 49.25 -
(Madry et al., 2019) + clat 51.64 -
WRN34-10 pgd-at 51.50 25.56
pgd-at + clat 52.88 27.62
(Croce & Hein, 2020) 57.3 -
(Croce & Hein, 2020) + clat 59.98 -
VGG19 pgd-at 40.42 19.54
pgd-at + clat 41.72 20.45
RN18 pgd-at 40.48 20.21
pgd-at + clat 42.86 21.76
twins (Liu et al., 2023) 47.89 25.45
twins + clat 51.39 28.12
autolora (Zhang et al., 2024) 48.95 27.48
autolora + clat 53.21 30.49
@@
Table 3 and Table 4 evaluate the robustness against black-box attacks (Auto Attack and PGD-AT respectively) between models trained solely using PGD-AT and those augmented with CLAT. Attack settings are the same as those of the white-box attacks. As a sanity check, the accuracies under black-box attack surpass those observed under white-box scenarios, indicating that gradient masking does not appear in the CLAT model, and that the white-box robustness evaluation is valid. More significantly, models trained with CLAT consistently outperform those trained with PGD-AT, maintaining superior resilience in both black-box and white-box settings, regardless of the attack method or models employed.
Table 3: Comparative Analysis of Black-box Auto Attack (AA) Accuracy on CIFAR-10 and CIFAR-100. Model in each row is the attacker and each column the victim.
Network Method CIFAR-10 Adv. Acc. (%) CIFAR-100 Adv. Acc. (%)
DN121 RN50 VGG19 RN18 DN121 RN50 VGG19 RN18
DN121 pgd-at - 52.50 44.21 45.45 - 27.64 23.21 24.26
pgd-at + clat - 55.83 47.53 48.92 - 29.89 26.71 26.93
RN50 pgd-at 54.23 - 43.56 43.24 27.91 - 23.02 23.51
pgd-at + clat 56.72 - 46.21 47.01 30.11 - 26.34 25.86
VGG19 pgd-at 55.32 55.45 - 46.72 27.84 28.15 - 24.20
pgd-at + clat 59.83 59.72 - 49.31 30.20 29.79 - 26.55
RN18 pgd-at 53.21 51.73 43.21 - 29.31 26.75 22.91 -
pgd-at + clat 56.75 54.45 46.53 - 31.25 29.41 25.95 -